OpenStudy (anonymous):

how do you find the limit of an equation with a delta symbole like this.............???????????????

OpenStudy (anonymous):

epsilon-delta

OpenStudy (anonymous):

missing the question something like \[\lim_{\Delta x\to 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}\] ?

OpenStudy (anonymous):

hold on my computerws slow i putting it up now

OpenStudy (anonymous):

Do you mean formally proving a limit using delta-epsilon notation?

OpenStudy (anonymous):

\[\lim_{Deltax \rightarrow 0} 2(x+Deltax-2x)/Deltax\]

OpenStudy (anonymous):

use the definition of a limit

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{2(x+h)-2x}{h}\] is easier to write

OpenStudy (anonymous):

do the algebra first, get \[\frac{2(x+h)-2x}{h}=\frac{2x+2h-2x}{h}=\frac{2h}{h}=2\]

OpenStudy (anonymous):

you will recognize this as the slope of the line \(y=2x\) since you are computing the change in \(y\) over the change in \(x\) since this is a line, the slope is constant, and it is 2

OpenStudy (anonymous):

notice that when you get done the algebra, you get a constant, that does not depend on \(x\) that is because this is a line slope doesn't change

OpenStudy (anonymous):

yess its and intro to calc AP

OpenStudy (anonymous):

Think about what satellite said, and how you could apply to finding the slope of non-linear functions (like a parabola).

OpenStudy (anonymous):

ok then, would this one work the same way ??....\[\lim_{h \rightarrow 0} (x+h)^{2}-2x/h\]

OpenStudy (anonymous):

\[x ^{2}+h ^{2}-x ^{2}/h\]

OpenStudy (anonymous):

I think you mean \[\lim_{h\rightarrow 0} \frac{(x+h)^2 - x^2}{h}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What does that evaluate to?

OpenStudy (phi):

use FOIL to do (x+h)(x+h)

OpenStudy (anonymous):

ok were do i go from here

OpenStudy (anonymous):

\[\lim_{h\rightarrow 0}\frac{(x+h)^2-x^2}{h}\\ \lim_{h\rightarrow 0}\frac{x^2+2xh+h^2-x^2}{h} \] Can you find the limit?

OpenStudy (anonymous):

do i know sub in 0 for h

OpenStudy (anonymous):

now not know

OpenStudy (phi):

always simplify first

OpenStudy (anonymous):

how much further can i simplify

OpenStudy (phi):

x^2-x^2 looks promising

OpenStudy (anonymous):

do i cross out the h's

OpenStudy (anonymous):

Yes, you can do that after removing the x^2's. Can you then plug in h=0 without dividing by 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

arent u left with 2x

OpenStudy (anonymous):

or did i do something wrong?

OpenStudy (anonymous):

Sorry for the late response (computer issues). Yes, that's correct. You get 2x. So if that limit gives you the slope, what does 2x mean?

OpenStudy (anonymous):

2/1 ? idk @mboorstin

OpenStudy (anonymous):

I'm talking about the fact that a parabola's slope is 2x. What that means is that you can find the slope of a non-linear function, but that the slope is variable depending on the location of the function. Not a big deal if you don't get it right now, you'll spend a few days in class talking about it. Your limits were correct, but that's why (f(x+h)-f(x))/h is so important.

OpenStudy (anonymous):

oh ok thanks you for the exlaination